# Panoptic: True Joint mmWave Communication and Sensing with Compressive Sidelobe Forming

### v1

@@ Abstract
The state-of-the-art integrated communication and sensing techniques separate the two functionalities either in time or frequency, or exploit two distinct hardware (e.g., mmWave radar and communication module) [11, 12, 13]. Therefore, despite promising prospects, these efforts do not achieve true joint communication and sensing functionalities111We refer to sensing as finding and tracking surrounding reflecting objects in real-time, which for example can form back-up paths in case of blockage. that share time, frequency, and hardware resources.
In this paper, we present Panoptic 222In Greek mythology, Argus Panoptes is a giant with 100 eyes, whose name means ”all-seeing”., a novel framework for integrated communication and sensing at the same time, and frequency band. Panoptic does not require additional hardware deployment (e.g., radar) and can be implemented using pervasive and widely available single-RF chain mmWave antenna arrays. The fundamental challenge behind realizing this goal is two-fold: (i) the communication and sensing functions have inherently different requirements, i.e., sensing needs time-varying directional beam scanning while communication demands a stable and accurately aligned directional beam to achieve high SNR; (ii) communication includes detection of unknown data symbols while sensing involves analyzing the properties of known signals (e.g., cross-correlation of received and known preambles) to extract the key components of the wireless medium.
Panoptic tackles these challenges by introducing a novel antenna modulation and compressive signal processing framework that capitalizes on channel sparsity. Specifically, our key idea is to maintain a directional mainlobe toward the communication node while simultaneously generating pseudorandom sidelobe perturbations to allow for reflection sensing in non-mainlobe directions. The high-level overview of Panoptic is depicted in Fig. 1 illustrating a bistatic integrated communication and sensing architecture. A mmWave data-carrying link is formed between the two communication parties while, at the same time, sidelobe perturbations on one end assist with environmental tracking. Specifically, a reflected path yields random albeit unique power and phase fluctuations at the receiver for a given suite of randomly shaped sidelobes. Such variations may act as a signature for path sensing when properly extracted from the overall superposed signals containing unknown complex-valued symbols.
@@
While our compressive sidelobe forming framework addresses challenge (i) explained above, the problem of joint data recovery and path sensing has remained unsolved. Specifically, the received signal is a combination of an unknown data-modulated signal that travels through a set of unknown paths. Further, the non-line-of-sight (NLOS) signatures are inherently weaker due to additional path loss and significantly smaller directivity gain along the sidelobe directions where such paths are perceived. Hence, the NLOS fingerprints may be masked under a much-stronger mainlobe signal. To tackle this, we mathematically model the received signal and present a new signal processing framework that performs consecutive symbol detection and then NLOS angle estimation by integrating information across several compressive sidelobe patterns. Therefore, Panoptic enables real-time and continuous reflector identification and tracking at zero additional overhead, i.e., through joint sensing and data reception in mobile environments. Indeed, such environmental inputs are the key to immediate link recovery under blockage, and realizing reliable and resilient mmWave networking.
@@
In addition, we have also experimentally investigated the tradeoff between communication and sensing functionalities. Specifically, by randomly turning off a larger subset of antennas, the sidelobe randomness space grows and better sensing accuracy is observed, albeit at the cost of a higher SNR reduction for the mainlobe communication link. Fortunately, this penalty alleviates when large antenna arrays are employed. Indeed, the length of the NLOS perturbation signature correlates with , where is the number of OFF antennas out of total antennas while the mainlobe directivity scales by a factor of . Interestingly, our results reveal that randomly turning off only 2 out of 64 antenna elements creates sufficiently large random space for accurate NLOS angle estimation (error of less than 1.5 degrees) while penalizing the communication SNR by only a negligible amount of 0.28 dB (which does not often result in downgrading the Modulation and Coding Scheme (MCS) of the communication link). Further, thanks to its compressive sensing nature, the number of required beam configurations only grows by with the number of antennas; therefore, Panoptic is scalable to future massive mmWave arrays. Finally, Panoptic is compliant with IEEE 802.11ay and 5G NR as it does not impose any changes to the communication protocol and frame structure.
@@ II-A Principles of Joint Communication and Sensing
Therefore, most existing efforts in JCS leverage known preambles and pilots that are already part of the IEEE 802.11 or 5G NR frame structures for sensing. Existing mmWave sensing systems have repurposed the beam training period in IEEE 802.11ad/ay, during which the AP adopts sequential directional beams toward different angles. By analyzing the received sector sweep frames, one can then track the channel multipath or the angular location of the transmitting or receiving node passively [15, 16, 17, 18, 19, 20]. However, the beam training period is a low-duty cycle event (e.g., every 100 ms) originally designed to find the best-aligned beams between the AP and stations. Similarly, in 5G NR protocol, the SSB bursts can be sent with a periodicity of 5, 10, 20, 40, 80, or 160 ms (where typically SSB bursts take 5 ms in duration) and leveraged for sensing [21, 9]. Clearly, a more frequent SSB burst can offer advantages for sensing at the cost of increased overhead (i.e., occupying channel resources that be otherwise used for data communication). We emphasize that, unfortunately, these techniques do not support truly joint communication and sensing: when the AP/BS is performing beam training, it can sense the environment but without data communication; when the AP and stations are communicating, sensing cannot be achieved.
@@ II-C Multi-Beam Full Duplex JCS Systems
Different from the above strategies, another line of work exploits spatial division between sensing and communication via beam manipulation [28, 29]. Although beamforming for communication and sensing separately has been studied extensively [30], its application in JCS comes with a different set of challenges. Specifically, communication and sensing have different beam requirements: sensing needs time-varying directional beam scanning while communication demands stable and accurately aligned directional beam to achieve high SNR. More recently, multi-beam designs using full-duplex radios have been proposed, where there is a fixed directional beam for downlink communication and the other sweeping beam for collecting the reflected signal back for sensing [31, 32, 33]. In such case, complex beamforming optimization is needed to mitigate the self-interference between communication and sensing at the full-duplex node [34, 35, 36]. In such full-duplex monostatic architectures, the transmitted signals are known at the co-located receiver and can be used as a reference to easily extract the desired reflection information from nearby objects. Overall, full duplex radios are not common in practical infrastructure settings, especially for mobile devices because of the increased hardware complexity and power consumption.
While our work falls into the general category of beam manipulation for JCS, Panoptic has a fundamentally different system model exploiting half-duplex single-chain analog phased arrays commonly available in all mmWave nodes. Further, the bistatic JCS architecture in Panoptic means that nodes might not be tightly synchronized posing additional challenges for sensing. Further, unlike multi-beam full duplex JCS systems that assume known data preambles for sensing, Panoptic is the first to jointly decode unknown data symbols and sense channel components at the same time.
@@ III-B Compressive Sidelobe Beam Design
We consider a sensing interval of seconds during which the channel variation is small enough to allow for a quasi-stationary assumption of channel paths and end users. Note that can be adapted as a control knob in Panoptic’s framework. During this sensing interval, the mmWave end nodes adopt different beam configurations. We can write the received signal for the symbol under the beam configuration as follows:
where and are the receiver and transmitter weight vectors, respectively, and is the zero-mean noise. The wireless channel consists of multipath components described by their AoA/AoD pair and complex channel gain . Hence, we can write as follows:
@@
In principle, the pseudorandom sidelobe perturbations can be realized in both TX arrays and RX arrays. However, in practice, changing the radiation profile on both sides at the same time may cause challenges in sensing and inference. This is because the weight vector variation on both ends (i.e., and in Eq. (2)) should be somewhat synchronized. In other words, the receiver should know the weight vector (i.e., sidelobe patterns) adopted at the TX for each signal reception. Hence, without loss of generality, we consider a fixed transmit beam and a varying receive radiation pattern. We highlight that the Panoptic’s framework can be directly applied to a fixed received beam and varying transmit sidelobe beams. In Sec. III-C3, we will illustrate the communication-sensing tradeoffs and discuss that it is desirable to implement Panoptic’s compressive sidelobes on the mmWave end node (TX or RX) with more antennas.
@@ III-C1 Symbol Detection
The key insight behind data detection in Panoptic is that the overall received signal contains a strong contribution from the mainlobe with a stable amplitude and phase profile that does not significantly vary with sidelobe perturbations. In other words, the much weaker sidelobe signals bonding off reflectors and reaching the receiver would have much lower power than the mainlobe signal. Specifically, by taking into consideration Eq. (2) and Eq. (5) and assuming that the TX beam is aligned to the RX, we can write
Given the random perturbations in the sidelobe, the NLOS components will add up randomly at the receiver while the mainlobe direction provides a fixed and high directivity gain. Hence, the received constellation does not deviate too much from the nominal constellation realized when all received are ON. Fig. 3 illustrates the received constellation before and after applying compressive sidelobe patterns. We can observe that the received symbols have slightly lower power and random yet small phase fluctuations. Nevertheless, a scaled decision boundary can still be used for symbol detection. Since the antenna subset size () is known at the receiver, finding the scaled decision boundary is straightforward by multiplying the original scaled boundary and .
We emphasize that the communication SNR can be modeled as SNR = , thus, the received SNR increases with the antenna subset size (i.e., the number of ON antennas). Hence there is a tradeoff between accurate symbol extraction and sensing performance: decreasing the antenna subset size hinders the communication SNR and may yield a higher symbol error rate but it will increase the random space for sidelobe sensing. We will explore it in Sec. III-C3.
@@ III-C2 Reflector Sensing Framework
Coherent vs Non-Coherent Sensing. In principle, minute perturbations in power and phase can both be leveraged for NLOS sensing. Extending the mathematical framework of Panoptic to include complex signatures is straightforward. Specifically, only slight changes to Eq. (7) to Eq. (9) are needed, i.e., we can write them as and . However, for practical considerations, non-coherent sensing has certain advantages over its coherent counterpart: first, it is more resilient to the phase noise, carrier frequency offset, and synchronization non-idealities commonly seen in commercial mmWave arrays [37] and second, it allows for aggregation of information across several packets that might not be tightly synchronized in phase. We emphasize that if accurate phase measurements are available, Panoptic can be easily extended to include phase readings in the AoA extraction framework, as discussed above. Finally, while the above modeling considers a narrowband channel, the extension of Panoptic to wideband channels with multiple subcarriers (i.e., OFDM) is straightforward. We will experimentally evaluate Panoptic in both coherent and non-coherent cases (under OFDM) in Sec. V-E.
@@
Specifically, the idea of compressive sidelobe beams originates from compressive sensing in which several pseudorandom weight vectors are used for AoA extraction [39]. Specifically, in compressive sensing, , where is i.i.d. and follows a uniform distribution over to ensure , and . This results in completely random radiation patterns that unfortunately fail to provide sufficient directivity gain for a communication link. Instead, in Panoptic, by maintaining a high-gain mainlobe and perturbing the sidelobes, we achieve the best of two worlds: high-gain communication link from the mainlobe and compressive AoA estimation in sidelobe directions. Previous works rigorously proved that the path estimation task requires unique patterns [39]. Intuitively, by removing the impact of the mainlobe (see Eq. (8)), the residual pattern resembles the pseudorandom compressive beams from the literature; hence, the number of measurements/configurations scales with the same log factor. We skip the detailed proof due to space limitations. This indicates that Panoptic is scalable to future massive mmWave MIMO settings.
@@ III-C3 Tradeoff between Communication and Sensing
In Panoptic, the communication and sensing functions share the same front-end, time, frequency, and even the same waveform. Here, we elaborate on the tradeoffs between these two functionalities. In principle, the communication SNR depends on the directivity gain along the mainlobe direction, and the AoA resolution depends on the level of unique perturbations created along sidelobes angles. Given a fixed antenna array size, the mainlobe strength is directly related to the number of ON antennas (). The random space for sidelobe perturbations, however, depends on . Therefore, an antenna subset size below would not benefit communication or sensing. When increases (), the communication SNR improves but the random space for sensing shrinks, and vice versa. It is also evident that implementing Panoptic on a mmWave node with a larger antenna array size (larger ) is desired as it can create the same sidelobe random space while turning off fewer antennas, imposing less penalty on the communication SNR. Hence, if there is an asymmetry between the array size at the two mmWave end nodes, running Panoptic on the node that has a larger array is advantageous. Interestingly, our evaluations suggest that randomly turning off only 2 antennas in a 64-element array creates sufficiently large random space for angular sensing of 1.5 degrees while imposing an SNR loss of less than 0.3 dB for the directional communication link.

### v2

@@ Abstract
The state-of-the-art integrated communication and sensing techniques separate the two functionalities either in time or frequency, or exploit two distinct hardware (e.g., mmWave radar and communication module) [11, 12, 13, 14]. Therefore, despite promising prospects, these efforts do not achieve true joint communication and sensing functionalities111We refer to sensing as finding and tracking surrounding reflecting objects in real-time, which for example can form back-up paths in case of blockage. that share time, frequency, and hardware resources.
In this paper, we present Panoptic 222In Greek mythology, Argus Panoptes is a giant with 100 eyes, whose name means “all-seeing”., a novel framework for integrated communication and sensing at the same time, and frequency band. Panoptic does not require additional hardware deployment (e.g., radar) and can be implemented using pervasive and widely available single-RF chain mmWave antenna arrays. The fundamental challenge behind realizing this goal is two-fold: (i) the communication and sensing functions have inherently different requirements, i.e., sensing needs time-varying directional beam scanning while communication demands a stable and accurately aligned directional beam to achieve high SNR; (ii) communication includes detection of unknown data symbols while sensing involves analyzing the properties of known signals (e.g., cross-correlation of received and known preambles) to extract the key components of the wireless medium.
Panoptic tackles these challenges by introducing a novel antenna modulation and compressive signal processing framework that capitalizes on channel sparsity. Specifically, our key idea is to maintain a directional mainlobe toward the communication node while simultaneously generating pseudorandom sidelobe perturbations to allow for reflection sensing in non-mainlobe directions. The high-level overview of Panoptic is depicted in Fig. 1 illustrating a bistatic integrated communication and sensing architecture. A mmWave data-carrying link is formed between the two communication parties while, at the same time, sidelobe perturbations on one end assist with environmental tracking. Specifically, a reflected path yields random, albeit unique power and phase fluctuations at the receiver for a given suite of randomly shaped sidelobes. Such variations may act as a signature for path sensing when properly extracted from the overall superposed signals containing unknown complex-valued symbols.
@@
While our compressive sidelobe forming framework addresses challenge (i) explained above, the problem of joint data recovery and path sensing has remained unsolved. Specifically, the received signal is a combination of an unknown data-modulated signal that travels through a set of unknown paths. Further, the non-line-of-sight (NLOS) signatures are inherently weaker due to additional path loss and significantly smaller directivity gain along the sidelobe directions where such paths are perceived. Hence, the NLOS fingerprints may be masked under a much stronger mainlobe signal. To tackle this, we mathematically model the received signal and present a new signal processing framework that performs consecutive symbol detection and then NLOS angle estimation by integrating information across several compressive sidelobe patterns. Therefore, Panoptic enables real-time and continuous reflector identification and tracking at zero additional overhead, i.e., through joint sensing and data reception in mobile environments. Indeed, such environmental inputs are the key to immediate link recovery under blockage, and realizing reliable and resilient mmWave networking.
@@
In addition, we have also experimentally investigated the tradeoff between communication and sensing functionalities. Specifically, by randomly turning off a larger subset of antennas, the sidelobe randomness space grows and better sensing accuracy is observed, albeit at the cost of a higher SNR reduction for the mainlobe communication link. Fortunately, this penalty alleviates when large antenna arrays are employed. Indeed, the length of the NLOS perturbation signature correlates with , where is the number of OFF antennas out of the total antennas while the mainlobe directivity scales by a factor of . Interestingly, our results reveal that randomly turning off only 2 out of 64 antenna elements creates sufficiently large random space for accurate NLOS angle estimation (error of less than 1.5 degrees) while penalizing the communication SNR by only a negligible amount of 0.28 dB (which does not often result in downgrading the Modulation and Coding Scheme (MCS) of the communication link). Further, thanks to its compressive sensing nature, the number of required beam configurations only grows by with the number of antennas; therefore, Panoptic is scalable to future massive mmWave arrays. Finally, Panoptic is compliant with IEEE 802.11ay and 5G NR as it does not impose any changes to the communication protocol and frame structure.
@@ II-A Principles of Joint Communication and Sensing
Therefore, most existing efforts in JCS leverage known preambles and pilots that are already part of the IEEE 802.11 or 5G NR frame structures for sensing. Existing mmWave sensing systems have repurposed the beam training period in IEEE 802.11ad/ay, during which the AP adopts sequential directional beams toward different angles. By analyzing the received sector sweep frames, one can then track the channel multipath or the angular location of the transmitting or receiving node passively [16, 17, 18, 19, 20, 21, 22]. However, the beam training period is a low-duty cycle event (e.g., every 100 ms) originally designed to find the best-aligned beams between the AP and stations. Similarly, in 5G NR protocol, the SSB bursts can be sent with a periodicity of 5, 10, 20, 40, 80, or 160 ms (where typically SSB bursts take 5 ms in duration) and leveraged for sensing [23, 9]. Clearly, a more frequent SSB burst can offer advantages for sensing at the cost of increased overhead (i.e., occupying channel resources that be otherwise used for data communication). We emphasize that, unfortunately, these techniques do not support truly joint communication and sensing: when the AP/BS is performing beam training, it can sense the environment but without data communication; when the AP and stations are communicating, sensing cannot be achieved.
@@ II-C Multi-Beam Full Duplex JCS Systems
Different from the above strategies, another line of work exploits spatial division between sensing and communication via beam manipulation [30, 31]. Although beamforming for communication and sensing separately has been studied extensively [32], its application in JCS comes with a different set of challenges. Specifically, communication and sensing have different beam requirements: sensing needs time-varying directional beam scanning while communication demands stable and accurately aligned directional beam to achieve high SNR. More recently, multi-beam designs using full-duplex radios have been proposed, where there is a fixed directional beam for downlink communication and the other sweeping beam for collecting the reflected signal back for sensing [33, 34, 35]. In such case, complex beamforming optimization is needed to mitigate the self-interference between communication and sensing at the full-duplex node [36, 37, 38]. In such full-duplex monostatic architectures, the transmitted signals are known at the co-located receiver and can be used as a reference to easily extract the desired reflection information from nearby objects. Overall, full duplex radios are not common in practical infrastructure settings, especially for mobile devices, because of the increased hardware complexity and power consumption.
In contrast to the prior art, Panoptic has fundamentally different hardware requirements and system goals: (i) Panoptic employs half-duplex single-chain analog phased arrays that are commonly available in all mmWave nodes. Hence, there is no hardware barrier for deploying Panoptic on existing infrastructure. (ii) Panoptic has a bistatic JCS architecture. It eliminates self-interference since transmission and reception occur at separate nodes. This makes it well-suited for deployment on commercially available, single-RF-chain mmWave arrays. Additionally, they align naturally with existing mobile network roles, supporting practical integration into real-world systems. Further, the bistatic JCS architecture in Panoptic means that nodes might not be tightly synchronized posing additional challenges for sensing. (iii) Finally, Panoptic aims to extract the unknown data symbols jointly with environmental components. This is not a problem in full-duplex settings, as the data is already known at the co-located TX/RX node. Panoptic is the first to jointly decode unknown data symbols and sense channel components at the same time.
@@ III-B Compressive Sidelobe Beam Design
We consider a sensing interval of seconds during which the channel variation is small enough to allow for a quasi-stationary assumption of channel paths and end users. Note that can be adapted as a control knob in Panoptic’s framework. During this sensing interval, the mmWave end nodes adopt different beam configurations , where the RX collects symbols under each beam configuration. We can write the received signal for the symbol under the beam configuration as follows:
where and are the receiver and transmitter weight vectors, respectively, is the symbol for each time slot, and is the zero-mean noise. The wireless channel consists of multipath components described by their AoA/AoD pair and complex channel gain . Hence, we can write as follows:
@@
In principle, the pseudorandom sidelobe perturbations can be realized in both TX arrays and RX arrays. However, in practice, changing the radiation profile on both sides at the same time may cause challenges in sensing and inference. This is because the weight vector variation on both ends (i.e., and in Eq. (2)) should be somewhat synchronized. In other words, the receiver should know the weight vector (i.e., sidelobe patterns) adopted at the TX for each signal reception. Hence, without loss of generality, we consider a fixed transmit beam and a varying receive radiation pattern.
We highlight that the Panoptic’s framework can also be directly applied to a fixed received beam and varying transmit sidelobe beams. In Sec. III-C3, we will illustrate the communication-sensing tradeoffs and discuss that it is desirable to implement Panoptic’s compressive sidelobes on the mmWave end node with more antennas (whether TX or RX). Furthermore, the fixed beam on one end is pointed toward the intended communication party. There is an inherent trade-off in selecting its beamwidth: a narrower beam improves communication SNR but restricts the spatial extent of illumination, limiting the sensing capability. In contrast, a wider beam increases the likelihood of illuminating nearby reflectors, enhancing the quality of sensing information, albeit at the cost of reduced SNR. This trade-off is well understood in the community and is independent of the proposed Panoptic framework.
@@ III-C1 Symbol Detection
The key insight behind data detection in Panoptic is that the overall received signal contains a strong contribution from the mainlobe with a stable amplitude and phase profile that does not significantly vary with sidelobe perturbations. In other words, the much weaker sidelobe signals bouncing off reflectors and reaching the receiver would have much lower power than the mainlobe signal. For instance, based on direct array factor calculation for a 16-element ULA with 4 randomly deactivated antennas (12 active), steered to broadside, the minimum amount of mainlobe-to-sidelobe-difference is 9.6 dB. This delta is much more pronounced when the extra path loss and reflection loss of NLOS components are taken into account.
Hence, by taking into consideration Eq. (2) and Eq. (5) and assuming that the TX beam is aligned to the RX, we can write
Given the random perturbations in the sidelobe, the NLOS components will add up randomly at the receiver while the mainlobe direction provides a fixed and high directivity gain. Hence, the received constellation does not deviate too much from the nominal constellation realized when all received are ON. Fig. 3 illustrates the received constellation before and after applying compressive sidelobe patterns. We can observe that the received symbols have slightly lower power and random yet small phase fluctuations. Nevertheless, a scaled decision boundary can still be used for symbol detection. Since the antenna subset size () is known at the receiver, finding the scaled decision boundary is straightforward by multiplying the original scaled boundary and . We use a standard approach for channel estimation and symbol detection. Since the number of antennas OFF remains unchanged, the decision boundary for demodulation also stays the same. As a result, we perform data communication with the conventional method by using preambles at the start of each packet and apply channel equalization accordingly. .
We emphasize that the communication SNR can be modeled as SNR = , where is the average energy of symbols and is the noise of Rx, thus, the received SNR increases with the antenna subset size (i.e., the number of ON antennas). Hence there is a tradeoff between accurate symbol extraction and sensing performance: decreasing the antenna subset size hinders the communication SNR and may yield a higher symbol error rate but it will increase the random space for sidelobe sensing. We will explore it in Sec. III-C3.
@@ III-C2 Reflector Sensing Framework
Coherent vs Non-Coherent Sensing. In principle, minute perturbations in power and phase can both be leveraged for NLOS sensing. Extending the mathematical framework of Panoptic to include complex signatures is straightforward. Specifically, only slight changes to Eq. (7) to Eq. (9) are needed, i.e., we can write them as and . However, for practical considerations, non-coherent sensing has certain advantages over its coherent counterpart: first, it is more resilient to the phase noise, carrier frequency offset, and synchronization non-idealities commonly seen in commercial mmWave arrays [39] and second, it allows for aggregation of information across several packets that might not be tightly synchronized in phase. We emphasize that if accurate phase measurements are available, Panoptic can be easily extended to include phase readings in the AoA extraction framework, as discussed above. Inclusion of phase can pave the way for phase-based advanced sensing, like Doppler extraction and motion sensing. Finally, while the above modeling considers a narrowband channel, the extension of Panoptic to wideband channels with multiple subcarriers (i.e., OFDM) is straightforward. We will experimentally evaluate Panoptic in both coherent and non-coherent cases (under OFDM) in Sec. V-E.
@@
Specifically, the idea of compressive sidelobe beams originates from compressive sensing in which several pseudorandom weight vectors are used for AoA extraction [41]. Specifically, in compressive sensing, , where is i.i.d. and follows a uniform distribution over to ensure , and . This results in completely random radiation patterns that, unfortunately, fail to provide sufficient directivity gain for a communication link. Instead, in Panoptic, by maintaining a high-gain mainlobe and perturbing the sidelobes, we achieve the best of two worlds: high-gain communication link from the mainlobe and compressive AoA estimation in sidelobe directions. Previous works rigorously proved that the path estimation task requires unique patterns [41], where is the array size. Intuitively, by removing the impact of the mainlobe (see Eq. (8)), the residual pattern resembles the pseudorandom compressive beams from the literature; hence, the number of measurements/configurations scales with the same log factor. We skip the detailed proof due to space limitations. We note that a scanning sidelobe design—where a fixed mainlobe supports communication and a directional sidelobe sequentially scans across angles—incurs a higher overhead of , as it requires a separate beam configuration for each direction.
This indicates that Panoptic is scalable to future massive mmWave MIMO settings.
Computational Complexity of Sensing Framework. The sensing algorithm is lightweight, with correlation being the most complex operation. If we note the number of theoretical compressive beams we use as , the time complexity is , where is the number of compressive sidelobe configurations we use and is the number of Rx antennas. The computational complexity mainly comes from theoretical compressive beam simulation. If we can simulate and store them beforehand, the running time complexity is only . Therefore, it does not require extra computational resources and can be directly applied to current hardware.
@@ III-C3 Tradeoff between Communication and Sensing
In Panoptic, the communication and sensing functions share the same front-end, time, frequency, and even the same waveform. Here, we elaborate on the tradeoffs between these two functionalities. In principle, the communication SNR depends on the directivity gain along the mainlobe direction, and the AoA resolution depends on the level of unique perturbations created along sidelobes angles. Given a fixed antenna array size, the mainlobe strength is directly related to the number of ON antennas (). The random space for sidelobe perturbations, however, depends on , where is the number of antennas in the RX array. Therefore, an antenna subset size below would not benefit communication or sensing. When increases (), the communication SNR improves but the random space for sensing shrinks, and vice versa. It is also evident that implementing Panoptic on a mmWave node with a larger antenna array size (larger ) is desired as it can create the same sidelobe random space while turning off fewer antennas, imposing less penalty on the communication SNR. Hence, if there is an asymmetry between the array size at the two mmWave end nodes, running Panoptic on the node that has a larger array is advantageous. Interestingly, our evaluations suggest that randomly turning off only 2 antennas in a 64-element array creates sufficiently large random space for angular sensing of 1.5 degrees while imposing an SNR loss of less than 0.3 dB for the directional communication link.